### v1

# Rademacher’s Theorem for Calderon-Zygmund-type Spaces

Abstract
Rademacher’s Theorem is a classical result stating that a Lipschitz function on $\mathbb{R}^{d}$ possesses a total differential almost everywhere. This implies that if $f$ is a function defined on $\mathbb{R}^{d}$ belonging to the Calderón–Zygmund space $T^{\infty}_{1}(x)$ for every $x\in\mathbb{R}^{d}$, then $f\in t^{\infty}_{1}(x)$ for almost every $x\in\mathbb{R}^{d}$.
The main purpose of this paper is to extend this property to a broader functional framework. More precisely, we replace the Lipschitz condition by the assumption that $f$ belongs to a general weighted Calderón–Zygmund space $T^{p}_{\phi}(x)$ for $x\in E$, where $E$ is a measurable subset of $\mathbb{R}^{d}$ and $\phi$ is a weight with fractional indices. We then show that, under suitable assumptions, the function belongs to $t^{p}_{\phi}(x)$ for almost every $x\in E$.
The result can be obtained either by imposing additional hypotheses or by working within a suitably adapted $t^{p}_{\phi}$ space. Whenever relevant, we also provide counterexamples illustrating the sharpness and the limitations of the statement.
Keywords: Rademacher’s Theorem, Calderón-Zygmund Spaces, Whitney’s Extension Theorem, Boyd functions
1 Introduction
The $T^{p}_{u}$ spaces were initially introduced by Calderón and Zygmund [3] : given a point $x_{0}\in\mathbb{R}^{d}$, $p\in[1,\infty]$ and a number $u\geq-d/p$, $T^{p}_{u}(x_{0})$ denotes the class of functions $f$ in $L^{p}(\mathbb{R}^{d})$ for which there exists a polynomial $P$ of degree strictly less than $u$ with the property that
$r^{-d/p}\|f-P\|_{L^{p}(B(x_{0},r))}\leq C\,r^{u}\quad\forall r>0.$ (1)
If, in addition,
$r^{-d/p}\|f-P\|_{L^{p}(B(x_{0},r))}=o(r^{u})\text{ when }r\rightarrow 0^{+},$ (2)
then $f\in t^{p}_{u}(x_{0})$. They allow to determine the pointwise regularity of functions that are not locally bounded (see for example [6, 16, 9]). Indeed, since the spaces $T^{p}_{u}$ are decreasing with $u$, one may define the $p$-exponent of $f$ :
$h_{p}^{f}(x_{0})=\sup\{u\geq-d/p:f\in T^{p}_{u}(x_{0})\}.$
If $p=\infty$, $T^{p}_{u}(x_{0})$ is the classical pointwise Hölder space, denoted by $\Lambda^{u}(x_{0})$.
Rademacher’s Theorem is a classical result which states that a Lipschitz function on $\mathbb{R}^{d}$ has a total differential almost everywhere. This means that if $f$ is a function defined on $\mathbb{R}^{d}$ which belongs to the space $T^{\infty}_{1}(x)$ for every $x\in\mathbb{R}^{d}$, then $f\in t^{p}_{1}(x)$ for almost every $x\in\mathbb{R}^{d}$, where the polynomial involved in the definition of the spaces $t^{p}_{1}(x)$ has degree less than or equal to 1.
A. Calderón generalized the differentiability of Sobolev functions by showing that if $\Omega\subset\mathbb{R}^{d}$ is bounded and open, then every function in $W^{p}_{1}(\Omega)$ is differentiable almost everywhere for $p>d$ [5]. This result follows naturally from the Lebesgue differentiation theorem and the Sobolev embedding theorem, and Rademacher’s theorem appears as a special case, since every Lipschitz function on $\Omega$ lies in $W^{\infty}_{1}(\Omega)$. Rademacher’s Theorem also admits a broader formulation for Lipschitz mappings from Euclidean spaces into arbitrary metric spaces, where differentiability is defined in terms of metric differentials rather than classical derivatives [4].
The original idea here was to replace the Lipschitz condition in Rademacher’s Theorem by the assumption that the function belongs to $T^{p}_{\phi}(x)$ for $x\in E$, where $E$ is a measurable subset of $\mathbb{R}^{d}$, and to show that the function then belongs to $t^{p}_{\phi}(x)$ for almost all $x\in E$, where space $T^{p}_{\phi}$ (resp. $t^{p}_{\phi}$) is a generalized version of $T^{p}_{u}$ (resp. $t^{p}_{u}$) by replacing $r\mapsto r^{u}$ appearing in (1) (resp. (2)) by a function parameter (see section 2). This yields an $L^{p}$ version of Rademacher’s Theorem, and this has been established for the spaces $T^{p}_{u}$, with $u\in\mathbb{N}_{0}$ [17].
Here, $\mathcal{L}$ stands for the Lebesgue measure, $\mathbb{N}_{0}=\{0,1,2,3,$...\} and $\mathbb{N}=\{1,2,3,..\}$.
2 Calderón-Zygmund spaces with Function parameter
A crucial aspect in order to refine regularity spaces is the notion of Boyd functions [1, 2, 7, 8, 14, 10, 11], which serve as a replacement for the traditional functions $t\mapsto t^{\theta}$ typically used in these settings: a function $\phi:(0,\infty)\to(0,\infty)$ is a Boyd function if it is continuous, $\phi(1)=1$ and
$\overline{\phi}(t):=\sup_{s>0}\frac{\phi(st)}{\phi(s)}<\infty,\quad\forall t\in(0,\infty).$
The lower and upper Boyd indices of a Boyd function $\phi$ are then defined by
$\underline{b}(\phi):=\sup_{t<1}\frac{\log\bar{\phi}(t)}{\log t}=\lim_{t\to 0}\frac{\log\bar{\phi}(t)}{\log t}$
and
$\overline{b}(\phi):=\inf_{t>1}\frac{\log\bar{\phi}(t)}{\log t}=\lim_{t\to\infty}\frac{\log\bar{\phi}(t)}{\log t},$
respectively.
Definition 1.
Let $x_{0}\in\mathbb{R}^{d}$, $p\in[1,\infty]$ and $\phi\in\mathcal{B}$ such that $\underline{b}(\phi)>-\frac{d}{p}$. A function $f\in L^{p}(\mathbb{R}^{d})$ belongs to the space $T^{p}_{\phi}(x_{0})$ if there exists a polynomial $P$ with degree strictly smaller than $\underline{b}(\phi)$ and a constant $C>0$ such that
$r^{-d/p}\|f-P\|_{L^{p}(B(x_{0},r))}\leq C\,\phi(r)\quad\forall r>0.$ (3)
The condition $\underline{b}(\phi)>-\frac{d}{p}$ is here to ensure that the spaces are not degenerate. One can easily check that the polynomial appearing in (3) is unique. We set
$\Lambda^{\phi}(x_{0})=T^{\infty}_{\phi}(x_{0}).$
Let $f\in T^{p}_{\phi}(x_{0})$ and $P$ be the polynomial relating to $f$, then $P$ can be written as
$P(x)=\sum_{|\alpha|<\underline{b}(\phi)}\frac{D^{\alpha}P(x_{0})}{|\alpha|!}(x-x_{0})^{\alpha}.$
We set
$|f|_{T^{p}_{\phi}(x_{0})}=\sup_{r>0}\phi(r)^{-1}r^{-d/p}\|f-P\|_{L^{p}(B(x_{0},r))},$
and
$\|f\|_{T^{p}_{\phi}(x_{0})}=\|f\|_{L^{p}(\mathbb{R}^{d})}+\sum_{|\alpha|<\underline{b}(\phi)}\frac{|D^{\alpha}P(x_{0})|}{|\alpha|!}+|f|_{T^{p}_{\phi}(x_{0})}.$
Then it is straightforward that $(T^{p}_{\phi}(x_{0}),\|\cdot\|_{T^{p}_{\phi}(x_{0})})$ is a Banach space. Let $t^{p}_{\phi,n}(x_{0})$, constitued of functions $f\in L^{p}(\mathbb{R}^{d})$ such that there exists a polynomial $P$ of degree at most $n$ such that
$r^{-d/p}\|f-P\|_{L^{p}(B(x_{0},r))}=o(\phi(r))\text{\emph{ when }}r\rightarrow 0^{+}.$
If $n<\underline{b}(\phi)$, it is easy to prove that $t^{p}_{\phi,n}(x_{0})$ is a closed subspace of $T^{p}_{\phi}(x_{0})$ with respect to the topology induced by the norm $\|\cdot\|_{T^{p}_{\phi}(x_{0})}$ (see [12]).
Remark 1.
If $\phi(t)=t^{u}$, with $u\in\mathbb{N}_{0}$, then $t^{p}_{\phi,u}(x_{0})$ is the usual space $t^{p}_{u}(x_{0})$ of [17]. Remark that, in this case, the space $T^{p}_{\phi}(x_{0})$ involves a polynomial of degree at most u1u-1 whereas the space $t^{p}_{u}(x_{0})$ involves a polynomial of degree at most $u$, so that $t^{p}_{u}(x_{0})$ is not a true subspace of $T^{p}_{\phi}(x_{0})$.
These spaces are used to capture finer notions of regularity. For example (see [15]), if $B$ is a Brownian motion on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, then almost surely and for almost every $t_{0}\in\mathbb{R}$, there exists a constant $C>0$ such that
$|B(t)-B(t_{0})|\leq C\big(|t-t_{0}|\log\log 1/|t-t_{0}|\big)^{1/2},$ (4)
whereas for every $t_{0}\in\mathbb{R}$, the Hölder exponent of $B$ at $t_{0}$ equals $h(t_{0})=\frac{1}{2}$. Therefore, if $\phi\in\mathcal{B}$ is such that $\phi(t)\sim(t\log\log 1/|t|)^{1/2}$ for all $t>0$, then $T^{\infty}_{\phi}$ is the adapted functional space that take into account the iterated logarithm in the pointwise regularity of $B$, a phenomenon that classical Hölder spaces fail to capture. In [12], they get pointwise estimates for solutions of elliptic partial differential equations in the context of $T^{p}_{\phi}$ spaces. They also get a more general version of Whitney’s extension Theorem.
Theorem 1.
(Generalization of Whitney’s Extension Theorem)
Let $E\subseteq\mathbb{R}^{d}$ be a closed set, $U=\{x\in\mathbb{R}^{d}\mid d(x,E)<1\}$, $p\in[1,\infty]$, $n\in\mathbb{N}_{0}$ and $\phi\in\mathcal{B}$ such that $n<\underline{b}(\phi)$.
[…]
$|\Delta^{\,m-n}_{h}D^{\alpha}F(x)|\leq C\,\phi(|h|)\,|h|^{-n}$
for every multi-index $\alpha$ with $|\alpha|=n$.
We generalize here to the spaces $T^{p}_{\phi}(x_{0})$ the approach carried out for spaces $T^{p}_{u}(x_{0})$ and $t^{p}_{u}(x_{0})$, $u\in\mathbb{N}_{0}$, in [17]. This generalization proceeds in two directions: the indices are allowed to be non-integer, and one may also consider more refined regularity spaces, with the flexibility that a weight can exhibit more nuanced behavior than a pure power function; for instance, involving iterated logarithmic terms.
3 Useful Results
The following Lemma is given in [17].
Lemma 1.
Let $n\in\mathbb{N}_{0}$. There exists a function $\varphi\in\mathcal{D}(\mathbb{R}^{d})$ with support contained in $\overline{B(0,1)}$ such that for every polynomial $P$ of degree less than or equal to $n$ and for every $\varepsilon>0$, we have $\varphi_{\varepsilon}\ast P=P$, where $\varphi_{\varepsilon}=\varepsilon^{-d}\varphi\left(\frac{\cdot}{\varepsilon}\right)$.
Lemma 2.
Let $E\subseteq\mathbb{R}^{d}$ be a measurable set, $p\in[1,\infty]$, and $\phi\in\mathcal{B}$ such that $\underline{b}(\phi)>-d/p$. If $f\in T^{p}_{\phi}(x_{0})$, then the function
$\xi_{f}:E\rightarrow[0,\infty):x\mapsto\|f\|_{T^{p}_{\phi}(x)}$
is measurable.
Proof.
Let $x\in E$. We denote by $P_{x}$ the polynomial associated with $f$ at $x$, of degree strictly less than $\underline{b}(\phi)$. It therefore suffices to show that for every multi-index $\alpha$ with $|\alpha|<\underline{b}(\phi)$,
$x\in E\mapsto D^{\alpha}P_{x}(x)$
is measurable. If $\underline{b}(\phi)\leq 0$, the claim is obvious since $P_{x}$ is the zero polynomial for every $x\in E$. We may thus assume that $\underline{b}(\phi)>0$.
Let $\varphi$ be the function from Lemma 1 for $n=\lfloor\underline{b}(\phi)\rfloor$.
We define, for every $x\in E$,
$R_{x}=f-P_{x}.$
For $\varepsilon>0$ and fixed $x\in E$, we have
$D^{\alpha}\left(\varphi_{\varepsilon}\ast f(x)\right)$ $=D^{\alpha}\left(\varphi_{\varepsilon}\ast P_{x}(x)\right)+D^{\alpha}\left(\varphi_{\varepsilon}\ast R_{x}(x)\right)$
$=D^{\alpha}P_{x}(x)+\int_{\mathbb{R}^{d}}\varepsilon^{-d-|\alpha|}(D^{\alpha}\varphi)\left(\frac{y}{\varepsilon}\right)R_{x}(x-y)\,dy.$
But, since $|\alpha|<\underline{b}(\phi)$, using properties of Boyd functions, we get
$\left|\int_{\mathbb{R}^{d}}\varepsilon^{-d-|\alpha|}\,(D^{\alpha}\varphi)\left(\frac{y}{\varepsilon}\right)\,R_{x}(x-y)\,dy\right|$ $\leq C_{\varphi}\,\varepsilon^{-d-|\alpha|}\,\int_{B(0,\varepsilon)}|R_{x}(x-y)|\,dy$
$=C_{\varphi}\,\varepsilon^{-|\alpha|}\,\varepsilon^{-d}\,\|f-P_{x}\|_{L^{1}(B(x,\varepsilon))}$
$\leq C_{\varphi}\,\varepsilon^{-|\alpha|}\,C_{d}^{1-1/p}\,\varepsilon^{-d/p}\,\|f-P_{x}\|_{L^{p}(B(x,\varepsilon))}$
$\leq C^{\prime}\,\varepsilon^{-|\alpha|}\,\phi(\varepsilon)$
$\leq C^{\prime\prime}\,\varepsilon^{-|\alpha|}\,\varepsilon^{\underline{b}(\phi)-((\underline{b}(\phi)-|\alpha|)/2)}$
$=C^{\prime\prime}\varepsilon^{(\underline{b}(\phi)-|\alpha|)/2)}\underset{\varepsilon\rightarrow 0^{+}}{\longrightarrow}0.$
Therefore, for all $x\in E$, $D^{\alpha}\left(\varphi_{\varepsilon}\ast f(x)\right)$ tends to $D^{\alpha}P_{x}(x)$, hence the conclusion. ∎
Lemma 3.
Let $p\in[1,\infty)$, $\phi\in\mathcal{B}$ such that $\underline{b}(\phi)>0$ and $f\in L^{p}(\mathbb{R}^{d})$, if
$r^{-d/p}\,\|f\|_{L^{p}(B(x,r))}\leq C\,\phi(r)\quad\forall r>0$
for all $x$ in a measurable set $E$ of $\mathbb{R}^{d}$, then for almost all $x\in E$,
$r^{-d/p}\|f\|_{L^{p}(B(x,r))}=o(\phi(r))\text{\emph{ when }}r\rightarrow 0^{+}.$
Proof.
We may suppose that $E$ is bounded and that $f$ is with compact support. Since $\underline{b}(\phi)>0$, there exists a constant $C_{\star}>0$ such that
$\overline{\phi}(r)\leq C_{\star}\,r^{\underline{b}(\phi)-\underline{b}(\phi)/2}\leq C_{\star}\quad\forall r\in(0,1].$
Let $\varepsilon>0$, we need to prove that for almost all $x\in E$,
$r^{-d/p}\|f\|_{L^{p}(B(x,r))}<\varepsilon\,\phi(r)$ (5)
for sufficiently small $r$.
Let $A$ be a closed subset of $E$ such that
$\mathcal{L}\,(E\backslash A)<\varepsilon.$
It is the sufficient to prove the result for almost all $x\in A$. For $x\in A$, one has
$\lim_{r\rightarrow 0^{+}}r^{-d}\,\|f\|_{L^{1}(B(x,r))}\leq C_{d}^{1-1/p}\lim_{r\rightarrow 0^{+}}r^{-d/p}\,\|f\|_{L^{p}(B(x,r))}\leq C^{\prime}\,\lim_{r\rightarrow 0^{+}}\phi(r)=0.$
Therefore, since almost all points are Lebesgue points, $f=0$ almost everywhere on $A$.
We define the open set $U$ by
$U=\{x\in\mathbb{R}^{d}:d(x,A)<1\}$
and for all $x\in U\backslash A$, we set
$h(x)=\frac{1}{10}\,d(x,A).$
Using classical covering Theorems (see for example Theorem 1.3.1 of [17]), there exists a countable set $S\subseteq U\backslash A$ such that the family
$\left\{\overline{B(s,h(s))}:s\in S\right\}$
is disjointed and
$\A.\bigcup_{s\in S}\overline{B(s,5h(s))}\supseteq U\backslash$ A.
We get
$\int_{A}\int_{U}\frac{|f(y)|}{|x-y|^{d}\,\phi(|x-y|)}dy\,dx$ $\leq\int_{A}\int_{U\backslash A}\frac{|f(y)|}{|x-y|^{d}\,\phi(|x-y|)}dy\,dx$
$\leq\int_{A}\sum_{s\in S}\int_{B(s,5h(s))}\frac{|f(y)|}{|x-y|^{d}\,\phi(|x-y|)}dy\,dx$
$=\sum_{s\in S}\int_{B(s,5h(s))}|f(y)|\int_{A}\frac{dx}{|x-y|^{d}\,\phi(|x-y|)}dy.$
Since $A$ is closed, for all $s\in S$, there exists $x_{s}\in A$ such that
$|s-x_{s}|=d(s,A)=10h(s),$
and
$B(s,5h(s))\subseteq B(x_{s},|s-x_{s}|+5h(s))=B(x_{s},15h(s)).$
Consequently,
$(5h(s))^{-d}\,\|f\|_{L^{1}(B(s,5h(s)))}$ $\leq C_{d}^{1-1/p}\,(5h(s))^{-d/p}\,\|f\|_{L^{p}(B(s,5h(s)))}$
$\leq C_{d}^{1-1/p}\,(5h(s))^{-d/p}\,\|f\|_{L^{p}(B(x_{s},15h(s)))}$
$\leq C^{\prime}\phi(15h(s))$
$\leq C^{\prime\prime}\phi(5h(s)).$
Remark that if $x\in A$ and $y\in B(s,5h(s))$, then
$|x-y|\geq d(s,A)-5h(s)=5h(s).$
Thus, for $y\in B(s,5h(s))$, we get
$\int_{A}\frac{dx}{|x-y|^{d}\,\phi(|x-y|)}$ $\leq C_{1}\int_{|5h(s)|}^{+\infty}\frac{r^{-1}}{\phi(r)}dr$
$\leq C_{1}\,C_{\star}\,(5h(s))^{\underline{b}(\phi)/2}\int_{|5h(s)|}^{+\infty}\frac{r^{-1}}{r^{\underline{b}(\phi)/2}\,\phi(5h(s))}\,dr$
$\leq C_{2}\,\phi(5h(s))^{-1}.$
We obtain
$\int_{A}\int_{B(s,5h(s))}\frac{|f(y)|}{|x-y|^{d}\,\phi(|x-y|)}dy\,dx$ $\leq C_{2}\,(\phi(5h(s))^{-1}\int_{B(s,5h(s))}|f(y)|\,dy$
$\leq C_{2}\,(\phi(5h(s))^{-1}\,(5(h(s))^{d}\,C^{\prime\prime}\,\phi(5h(s))$
$\leq C_{3}\,h(s)^{d}.$
Since family $\left\{\overline{B(s,h(s))}:s\in S\right\}$ is disjointed,
$\int_{A}\int_{U}\frac{|f(y)|}{|x-y|^{d}\,\phi(|x-y|)}dy\,dx\leq C_{3}\sum_{s\in S}h(s)^{d}\leq C_{3}\sum_{s\in S}h(s)<\infty,$
so that
$\int_{U}\frac{|f(y)|}{|x-y|^{d}\,\phi(|x-y|)}dy<\infty$
for almost all $x\in A$. Since $f$ has compact support,
$\int_{\mathbb{R}^{d}\backslash U}\frac{|f(y)|}{|x-y|^{d}\,\phi(|x-y|)}dy<\infty$
for all $x\in A$ and thus
$\int_{\mathbb{R}^{d}}\frac{|f(y)|}{|x-y|^{d}\,\phi(|x-y|)}dy<\infty$
for almost all $x\in A$.
We observe that, to establish the above result, we have merely used the fact that
$r^{-d}\,\|f\|_{L^{1}(B(x,r))}\leq C_{d}^{1-1/p}\,r^{-d/p}\,\|f\|_{L^{p}(B(x,r))}\leq C\phi(r).$
Therefore, if $g=|f|^{p}$, the assumption becomes
$r^{-d}\,\|g\|_{L^{1}(B(x,r))}\leq C\phi(r)^{p}$
for all $x\in E$.
But, if $\phi\in\mathcal{B}$, then $\phi^{p}\in\mathcal{B}$ et $\underline{b}(\phi^{p})=p\,\underline{b}(\phi)>0$.
Thus, by applying once again the argument used above, we obtain that
$\int_{\mathbb{R}^{d}}\frac{|f(y)|^{p}}{|x-y|^{d}\,\phi^{p}(|x-y|)}\,dy=\int_{\mathbb{R}^{d}}\frac{|g(y)|}{|x-y|^{d}\,\phi^{p}(|x-y|)}\,dy<+\infty$
for almost all $x\in E$.
For such $x$ and for all sufficiently small $r$, we have
$\int_{B(x,r)}\frac{|f(y)|^{p}}{|x-y|^{d}\,\phi^{p}(|x-y|)}<\varepsilon^{p}\,C_{\star}^{-p},$
and therefore, for all sufficiently small $r$,
$r^{-d/p}\|f\|_{L^{p}(B(x,r))}$ $=\phi(r)\left(\int_{B(x,r)}\frac{|f(y)|^{p}}{r^{d}\,\phi^{p}(r)}\,dy\right)^{1/p}$
$\leq\phi(r)\left(\int_{B(x,r)}\frac{|f(y)|^{p}}{r^{d}\,\phi^{p}(|x-y|)}\,dy\,C_{\star}^{p}\right)^{1/p}$
$<\varepsilon\,\phi(r).$
To generalize Rademacher’s theorem, we will also make use of Lusin’s Theorem, which establishes a connection between measurability and continuity for functions defined on locally compact Hausdorff spaces equipped with a Radon measure ([13]). We state a weaker version of this result for spaces $E$ contained in a Euclidean space $\mathbb{R}^{d}$, under the additional assumption that $E$ is bounded.
Theorem 2.
(Lusin’s Theorem)
A real-valued function $f$ defined on a bounded measurable subset $E$ of $\mathbb{R}^{d}$ is measurable if and only if, for every $\varepsilon>0$, there exists a closed set $E_{\varepsilon}\subseteq E$ such that
$\mathcal{L}(E\backslash E_{\varepsilon})\leq\varepsilon$
and such that $f$ is continuous on $E_{\varepsilon}$.
We also need some preliminary results concerning Sobolev spaces. Let $\Omega$ be an open set of $\mathbb{R}^{d}$.
Theorem 3.
If $f\in L^{p}(\Omega)$, then $f\in W^{p}_{1}(\Omega)$ if and only if there exists a function $F$ that coincides with $f$ almost everywhere on $\Omega$, which is absolutely continuous on almost every line segment of $\Omega$ axis-parallel and whose first-order partial derivatives belong to $L^{p}(\Omega)$.
The following result is also proved in [17] (see Theorems 3.4.1 and 3.4.2 of [17]) in a slightly more general setting where the null measure set in question is actually of zero Bessel capacity.
Theorem 4.
If $p>1$ and $1\leq m_{1}\leq m_{2}$ verify $(m_{2}-m_{1})\,p<d$ and $f\in W^{p}_{m_{2}}(\mathbb{R}^{d})$, then there exists a polynomial $Q$ of degree at most $m_{1}$ such that
$r^{-m_{1}}\,r^{-d/p}\,\left\|f-Q\right\|_{L^{p}(B(x,r))}\underset{r\rightarrow 0^{+}}{\longrightarrow}0$
for almost all $x\in\mathbb{R}^{d}$.
We are now in a position to generalize Rademacher’s Theorem. We make the proof of it [17] since we will need similar arguments for the new results.
Theorem 5.
(Rademacher’s Theorem in the $L^{p}$ context)
Let $E\subseteq\mathbb{R}^{d}$ measurable, $p\in(1,\infty)$ and $n\in\mathbb{N}_{0}$. If $f\in T^{p}_{n}(x)$ for all $x\in E$, then $f\in t^{p}_{n}(x)$ for almost all $x\in E$.
Proof.
$\bullet$ Let us first assume that $E$ is bounded. By lemma 2 and Lusin’s Theorem, there exist compacts $E_{1/m}\subseteq E$, $m\in\mathbb{N}$, such that
$\mathcal{L}\left(E\backslash E_{1/m}\right)\leq\frac{1}{m}$
and such that
$\xi_{f}:E_{1/m}\rightarrow[0,\infty):x\mapsto\|f\|_{T^{p}_{\phi}(x)}$
is continuous. Let $m\in\mathbb{N}$, there exists a constant $M=M(m)>0$ such that
$\|f\|_{T^{p}_{n}(x)}\leq M\quad\forall x\in E_{1/m}.$
Let $x\in E_{1/m}$, we denote by $P_{x}$ the polynomial of degree strictly less than or equal to n1n-1 associated with $f$. Thanks to Whitney’s extension Theorem 1, there exists an open set $U$ containing $E_{1/m}$ and a function $F\in C^{n-1}(U)$, with Lipchitz-continuous partial derivatives such that $F=f$ almost everywhere $E_{1/m}$ and such that for all $|\beta|<n$ and for almost all $x\in E_{1/m}$,
$D^{\beta}F(x)=D^{\beta}P_{x}(x).$
By Theorem 3, $F\in W^{p}_{n,\text{loc}}(\mathbb{R}^{d})$. Using Theorem 4 with $m_{1}=m_{2}=n$, it follows that for almost every $x\in\mathbb{R}^{d}$, there exists a polynomial $Q_{x}$ of degree at most $n$ such that
$r^{-d/p}\,\|F-Q_{x}\|_{L^{p}(B(x,r))}=o(r^{n})\quad\text{when }r\rightarrow 0^{+}.$
Moreover, for almost every $x\in E_{1/m}$,
$r^{-d/p}\,\|f-F\|_{L^{p}(B(x,r))}\leq C\,r^{n}.$
Using Lemma 3, we get
$r^{-d/p}\|f-F\|_{L^{p}(B(x,r))}=o(r^{n})\text{ when }r\rightarrow 0^{+},$
Thus, for almost all $x\in E_{1/m}$,
$r^{-d/p}\,\|f-Q_{x}\|_{L^{p}(B(x,r))}$ $\leq r^{-d/p}\,\|f-F\|_{L^{p}(B(x,r))}+r^{-d/p}\,\|F-Q_{x}\|_{L^{p}(B(x,r))}$
$=o(\phi(r))+o(r^{n+1})$
$=o(\phi(r))\text{ when }r\rightarrow 0^{+}.$
We have thus proved the result for sets $E_{1/m}$, $m\in\mathbb{N}$. Since a countable union of negligible sets is negligible, we deduce the conclusion for $E$.
$\bullet$ If $E$ is unbounded, it suffices to apply the previous case to the sets
$E^{(j)}=E\cap B(0,j),\quad j\in\mathbb{N},$
and to conclude again using the fact that a countable union of negligible sets is negligible. ∎
4 Rademacher’s Theorem for Calderon-Zygmund-type Spaces
We suppose that $\phi\in\mathcal{B}$ and $n\in\mathbb{N}_{0}$ are such that $n<\underline{b}(\phi)\leq\overline{b}(\phi)<n+1$. For those fractional indices, we can’t easily adapt the proof because of Theorem 4. The idea to get results is to reduce to the case of integer indices. Before the Theorem, let us remark that if we have some assumption on $f$ at a neighborhood of $x_{0}$, then $f\in t^{p}_{\phi,n+1}(x_{0})$.
Remark 2.
If $f\in C^{n+1}(V)$, where $V$ is an open neighborhood of $x_{0}$, then
$f\in t^{p}_{\phi,n}(x_{0})\subseteq t^{p}_{\phi,n+1}(x_{0}).$
Indeed, there exists $R>0$ such that $\overline{B(x_{0},R)}\subseteq V$. If $P$ denotes the Taylor polynomial of $f$ of order $n$ at $x_{0}$, then since $f$ is of class $C^{n+1}$ on $V$, there exists a constant $C>0$ such that
$|f(x)-P(x)|\leq C|x-x_{0}|^{n+1}\quad\forall x\in B(x_{0},R).$
Thus, for every $r\in(0,R]$, we have
$r^{-d/p}\|f-P\|_{L^{p}(B(x_{0},r))}\leq C^{\prime\prime}r^{n+1}.$
By assumption, there exists $\varepsilon>0$ such that $\overline{b}(\phi)+\varepsilon<n+1$. It follows that
$\phi(r)^{-1}r^{-d/p}\|f\|_{L^{p}(B(x_{0},r))}\leq\phi(r)^{-1}C^{\prime\prime}r^{n+1}\leq\frac{C^{\prime\prime}}{C_{1}}r^{\,n+1-(\overline{b}(\phi)+\varepsilon)}\underset{r\to 0^{+}}{\longrightarrow}0,$
and hence $f\in t^{p}_{\phi,n}(x_{0})$.
Remark that,
$t^{p}_{\overline{b}(\phi)}(x_{0})\subseteq t^{p}_{\phi,n}(x_{0})\subseteq t^{p}_{\phi,n+1}(x_{0})\subseteq t^{p}_{\underline{b}(\phi)}(x_{0}).$
Since the second inclusion is often strict, the best general Rademacher’s Theorem we can get is the following.
Theorem 6.
(Rademacher’s Theorem for Calderon-Zygmund-type Spaces)
Let $E\subseteq\mathbb{R}^{d}$ measurable, $p\in(1,\infty)$, $\phi\in\mathcal{B}$ and $n\in\mathbb{N}_{0}$ such that $n<\underline{b}(\phi)\leq\overline{b}(\phi)<n+1$. If $f\in T^{p}_{\phi}(x)$ for all $x\in E$, then $f\in t^{p}_{\phi,n+1}(x)$ for almost all $x\in E$.
Proof.
$\bullet$ Let us first assume that $E$ is bounded. By lemma 2 and Lusin’s Theorem, there exist compacts $E_{1/m}\subseteq E$, $m\in\mathbb{N}$, such that
$\mathcal{L}\left(E\backslash E_{1/m}\right)\leq\frac{1}{m}$
and such that
$\xi_{f}:E_{1/m}\rightarrow[0,\infty):x\mapsto\|f\|_{T^{p}_{\phi}(x)}$
is continuous. Let $m\in\mathbb{N}$, there exists a constant $M=M(m)>0$ such that
$\|f\|_{T^{p}_{\phi}(x)}\leq M\quad\forall x\in E_{1/m}.$
Let $x\in E_{1/m}$, we denote by $P_{x}$ the polynomial of degree strictly less than or equal to $n$ associated with $f$. Thanks to Whitney’s extension Theorem 1, there exists an open set $U$ containing $E_{1/m}$ and a function $F\in C^{n}(U)$ such that $F=f$ almost everywhere $E_{1/m}$ and such that for all $|\beta|\leq n$ and for almost all $x\in E_{1/m}$,
$D^{\beta}F(x)=D^{\beta}P_{x}(x).$
By Theorem 3, $F\in W^{p}_{n+1,\text{loc}}(\mathbb{R}^{d})$. Using Theorem 4 with $m_{1}=m_{2}=n+1$, it follows that for almost every $x\in\mathbb{R}^{d}$, there exists a polynomial $Q_{x}$ of degree at most $n$ such that
$r^{-d/p}\,\|F-Q_{x}\|_{L^{p}(B(x,r))}=o(r^{n+1})\quad\text{when }r\rightarrow 0^{+}.$
Moreover, for almost every $x\in E_{1/m}$,
$r^{-d/p}\,\|f-F\|_{L^{p}(B(x,r))}\leq C\,\phi(r).$
Using Lemma 3, we get
$r^{-d/p}\|f-F\|_{L^{p}(B(x,r))}=o(\phi(r))\text{ when }r\rightarrow 0^{+},$
Thus, for almost all $x\in E_{1/m}$,
$r^{-d/p}\,\|f-Q_{x}\|_{L^{p}(B(x,r))}$ $\leq r^{-d/p}\,\|f-F\|_{L^{p}(B(x,r))}+r^{-d/p}\,\|F-Q_{x}\|_{L^{p}(B(x,r))}$
$=o(\phi(r))+o(r^{n+1})$
$=o(\phi(r))\text{ when }r\rightarrow 0^{+}.$
We have thus proved the result for sets $E_{1/m}$, $m\in\mathbb{N}$. Since a countable union of negligible sets is negligible, we deduce the conclusion for $E$.
$\bullet$ If $E$ is unbounded, it suffices to apply the previous case to the sets
$E^{(j)}=E\cap B(0,j),\quad j\in\mathbb{N},$
and to conclude again using the fact that a countable union of negligible sets is negligible. ∎
Example 1.
Suppose that $\overline{b}(\phi)<h_{p}^{f}(x)$ for all $x\in E$, then the statement
$f\in T^{p}_{\phi}(x)\quad\forall x\in E\implies f\in t^{p}_{\phi,n}(x)\quad\text{for almost all }x\in E$
is always true since, in this case, $f\in t^{p}_{\phi,n}(x)$ for all $x\in E$.
Example 2.
The following observation with Brownian motion led us to conclude that Rademacher’s theorem does not hold for fractional indices in the space $t^{p}_{\phi,0}$ (as opposed to $t^{p}_{\phi,1}$). Indeed, using (4), it is clear that almost surely and for almost every $t_{0}\in\mathbb{R}$, $B\in T^{p}_{\phi}(t_{0})$, where $\phi(t)\sim(t\log\log 1/|t|)^{1/2}$ for all $t>0$ but almost surely and for almost every $t_{0}\in\mathbb{R}$,
$\limsup_{r\rightarrow 0^{+}}\frac{r^{-1/p}(\int_{B(t_{0},r)}|B(t)-B(t_{0})|^{p}dt)^{1/p}}{\sqrt{r\log\log 1/r}}=\sqrt{2}.$
5 Conclusion
Let $\phi\in\mathcal{B}$ and $n\in\mathbb{N}_{0}$ are such that $n<\underline{b}(\phi)\leq\overline{b}(\phi)<n+1$. On the one hand, we have the inclusions
$t^{p}_{\phi,n}(x_{0})\subseteq T^{p}_{\phi}(x_{0})\quad\text{and}\quad t^{p}_{\phi,n}(x_{0})\subseteq t^{p}_{\phi,n+1}(x_{0}),$
but in general we do not have the inclusion
$t^{p}_{\phi,n+1}(x_{0})\subseteq T^{p}_{\phi}(x_{0}).$
On the other hand, our Rademacher’s Theorem 6 states that
$f\in T^{p}_{\phi}(x)\,\,\,\,\,\,\forall x\in E\quad\implies\quad f\in t^{p}_{\phi,n+1}(x)\quad\text{for almost all}\ x\in E.$
Moreover, Example 1 shows that this latter result is only of pointwise interest when $\overline{b}(\phi)\geq h_{p}^{f}(x)$, since otherwise one already has $f\in t^{p}_{\phi,n}(x)$. Example 2 further illustrates the limitations of our result, and shows that in general one cannot expect $f$ to belong almost everywhere to $t^{p}_{\phi,n}(x)$.
Declarations
  • Funding : The author did not receive any specific funding for this work.
  • Ethics approval : Not applicable.
  • Availability of data and materials : Not applicable.